OpenStudy (anonymous):

Solve the equation 3/2 - 1/x-4 = -2/2x-8 . can someone explain how to solve this?

hero (hero):

If I had a quarter for every time I saw this question...

hero (hero):

\[\frac{3}{2} - \frac{1}{x-4} = -\frac{2}{2x-8}\]

hero (hero):

\[\frac{3}{2} - \frac{1}{x-4} = -\frac{2}{2(x-4)}\]

hero (hero):

\[\frac{3}{2} - \frac{1}{x-4} = -\frac{1}{x-4}\]

hero (hero):

\[\frac{3}{2} = \frac{1}{x-4} -\frac{1}{x-4}\]

hero (hero):

\[\frac{3}{2} = 0\] False, therefore no solution

hero (hero):

any questions?

OpenStudy (anonymous):

no thanks!

hero (hero):

Have any others like this?

OpenStudy (anonymous):

no, this is the only one